OpenStudy (anonymous):

I WILL GIVE MEDAL AND FAN: Find the center, vertices, and foci of the ellipse with equation ((x^2/)/(100)) + ((Y^2)/(64))= 1

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@ganeshie8

OpenStudy (usukidoll):

standard ellipse equation For a wider-than-tall ellipse with center at (h, k), having vertices a units to either side of the center and foci c units to either side of the center, the ellipse equation is: \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] For a taller-than-wide ellipse with center at (h, k), having vertices a units above and below the center and foci c units above and below the center, the ellipse equation is: \[\frac{(y-h)^2}{a^2}+\frac{(x-k)^2}{b^2}=1\] standard foci equation \[a^2-c^2=b^2\] standard center \[(h,k)\]

OpenStudy (usukidoll):

based on your equation \[\frac{x^2}{100}+\frac{y^2}{64}=1\] a > b there's a horizontal major axis

OpenStudy (usukidoll):

EXAMPLE: suppose we have the equation \[\frac{x^2}{4}+\frac{y^2}{1}=1\]

OpenStudy (usukidoll):

since a > b we have a horizontal major axis so we re-write this equation as where a = 2 and b = 1 \[\frac{x^2}{2^2}+\frac{y^2}{1^2}=1\]

OpenStudy (usukidoll):

before we start graphing, we want the center. Recall that the standard ellipse equation is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 \] For this equation we just have \[x^2 \] and \[y^2 \] \[\frac{x^2-0}{2^2}+\frac{y^2-0}{1^2}=1\] which means that our center \[(h,k) \rightarrow (0,0)\]

OpenStudy (usukidoll):



OpenStudy (usukidoll):

for vertices we need (-a,0), (a,0), (-b,0),(0,b) since a = 2 and b = 1 our vertex points are (-2,0) (2,0) (-1,0)(0,1)

OpenStudy (usukidoll):